# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ A.3 Detailed Balance, Prior Incorporation, and Stochasticity
still satisfy the Kolmogorov equation. The detailed balance condition ensures that the outflow, $p_{t|1}(z_{t}|z_{1})R^{\text{DB}}_{t}(z_{t},z_{t+\text{d}t}|z_{1})$, and inflow, $p_{t|1}(z_{t+\text{d}t}|z_{1})R^{\text{DB}}_{t}(z_{t+\text{d}t},z_{t}|z_{1})$, of probability mass to any given state are perfectly balanced. Under these conditions, this additive component’s contribution to the Kolmogorov equation becomes null (similar to the target guidance, as shown in the proof of of Lemma 10, in Sec. A.2).
@@
A.4 Hyperparameter Optimization Pipeline
A significant advantage of flow matching methods is their inherently greater flexibility in the sampling process compared to diffusion models, as they are more disentangled from the training stage. Each of the proposed optimization strategies exposed in Sec. 4.2.2 expands the search space for optimal performance. However, conducting a full grid search across all those methodologies is impractical for the computational resources available. To address this challenge, our sampling optimization pipeline consists of, for each of the proposed optimization strategies, all hyperparameters are held constant at their default values except for the parameter controlling the chosen strategy, over which we perform a sweep. The optimal values obtained for each strategy are combined to form the final configuration. In Tab. 6, we present the final hyperparameter values obtained for each dataset. This pipeline is sufficient to achieve state-of-the-art performance, which reinforces the expressivity of DeFoG. We expect to achieve even better results if a more comprehensive search of the hyperparameter space was carried out.
To better mode detailedly illustrate the influence of each sampling optimization, we show in Figures 8 and 10 in present the impact of varying parameter values across the synthetic datasets Figure 8 and molecular datasets in Figure 10 used in this work.
@@ A.4 Hyperparameter Optimization Pipeline
Figure 10: Sampling efficiency improvement over all molecular datasets. The Guacamol and MOSES datasets are evaluated with 2,000 samples, and validity and novelty are computed locally rather than using the original benchmark for efficiency.
While Figures 8 and 10 are highly condensed, we provide a more fine-grained version that specifically illustrates the influence of the hyperparameters $\eta$ and $\omega$. This version highlights their impact when generating with the full number of steps (500 and 1000 for molecular and synthetic data, respectively) and with 50 steps. As emphasized in Figures 11 and 12, the influence of these hyperparameters varies across datasets and exhibits distinct behaviors depending on the number of steps used.
@@
A.5 Understanding the Sampling Dynamics of $R^{*}_{t}$
In this section, we aim to provide deeper intuition into the sampling dynamics imposed by the design of $R^{*}_{t}$, as proposed by Campbell et al. (2024). The explicit formulation of $R^{*}_{t}$ can be found in Eq. 3. Notably, the denominator in the expression serves as a normalizing factor, meaning the dynamics of each sampling step are primarily influenced by the values in the numerator. Specifically, we observe the following relationship:
@@ A.5 Understanding the Sampling Dynamics of $R^{*}_{t}$
Table 4: Values of $R^{*}_{t}$ for different $z_{t}$ and $z_{t+\mathrm{d}t}$.
Condition $R^{*}(x_{t},j|z_{1})$ Intuition
@@
From the first two lines of Tab. 4, we observe that once the system reaches the predicted state $z_{1}$, it remains there. If not, $R^{*}_{t}$ only encourages transitions to other states under two conditions: either the target state is $z_{1}$ (third line), or the corresponding entries in the initial distribution for potential next states have smaller values than the current state (fourth line). As a result, the sampling dynamics are heavily influenced by the initial distribution, as discussed further in Sec. B.1.
@@
For instance, with the masking distribution, the fourth line facilitates transitions to states other than the virtual “mask” state, whereas for the uniform distribution, no transitions are allowed. For the marginal distribution, transitions are directed toward less likely states. Note that while these behaviors hold when the rate matrix consists solely of $R^{*}_{t}$, additional transitions can be introduced through $R^{\mathrm{DB}}_{t}$ (as detailed in Sec. A.3) or by applying target guidance (see Sec. A.2).
A.6 Performance Improvement for Undertrained Models
@@
In this section, we present the performance of a model trained on the QM9 dataset and the Planar dataset using only $30\%$ of the epochs compared to the final model being reported. We employ the same hyperparameters with Tab. 8 and Tab. 1 for the sampling setup, as reported in Tab. 6.
@@
Compared to fully trained models, our model achieves 99.0 Validity (vs. 99.3) and 96.4 Uniqueness (vs. 96.3) on the QM9 dataset. For the Planar dataset, it attains 95.5 Validity (vs. 99.5) and an average ratio of 1.4 (vs. 1.6). These results demonstrate that, even with significantly fewer training epochs, the model maintains competitive performance under a well-designed sampling procedure, although extended training can still further improve performance. Notably, all metrics surpass the discrete-time diffusion benchmark DiGress (Vignac et al., 2022). As a result, the optimization in the sampling stage proves particularly beneficial when computational resources are limited, by enhancing the performance of an undertrained model.
Appendix B Train Optimization
@@
In this section, we provide a more detailed analysis of the influence of the various training optimization strategies introduced in Sec. 4.2.1. In Sec. B.1, we empirically demonstrate the impact of selecting different initial distributions on performance, while in Sec. B.2, we examine the interaction between training and sampling optimization.
B.1 Initial Distributions
@@
Under DeFoG’s framework, the noising process for each dimension is modeled as a linear interpolation between the clean data distribution (the one-hot representation of the current state) and an initial distribution, $p_{0}$. As such, it is intuitive that different initial distributions result in varying performances, depending on the denoising dynamics they induce. In particular, they have a direct impact on the sampling dynamics through $R^{*}_{t}$ (see Sec. A.5) and may also pose tasks of varying difficulty for the graph transformer. In this paper, we explore four distinct initial distributions [^2] Recall that $Z$ represents the cardinality of the state space, and $\Delta^{Z-1}$ the associated probability simplex.:
@@
Uniform: $p_{0}=\left[\frac{1}{Z},\frac{1}{Z},\ldots,\frac{1}{Z}\right]\in\Delta^{Z-1}$. Here, the probability mass is uniformly distributed across all states, as proposed by Campbell et al. (2024).
@@
Masking: $p_{0}=[0,0,\ldots,0,1]\in\Delta^{Z}$. In this setting, all the probability mass collapses into a new “mask” state at $t=0$, as introduced by Campbell et al. (2024).
@@
Marginal: $p_{0}=[m_{1},m_{2},\ldots,m_{Z}]\in\Delta^{Z-1}$, where $m_{i}$ denotes the marginal probability of the $i$-th state in the dataset. This approach is widely used in state-of-the-art graph generation models (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024).
@@
Absorbing: $p_{0}=[0,\ldots,1,\ldots,0]\in\Delta^{Z-1}$, representing a one-hot encoding of the most common state (akin to applying an $\operatorname{argmax}$ operator to the marginal initial distribution).
@@
In Figure 17, we present the training curves for each initial distribution for three different datasets.
@@
We observe that the marginal distribution consistently achieves at least as good performance as the other initial distributions. This, along with the theoretical reasons outlined in Sec. 4.2.1, reinforces its use as the default initial distribution for DeFoG. The only dataset where marginal was surpassed was the SBM dataset, which we attribute to its inherently different nature (stochastic vs. deterministic). In this case, the absorbing distribution emerged as the best-performing choice. Interestingly, the absorbing distribution also tends to converge faster across datasets.
@@
B.2 Interaction between Sample and Train Distortions
From Sec. A.4, we observe that time distortions applied during the sampling stage can significantly affect performance. This suggests that graph discrete flow models do not behave evenly across time and are more sensitive to specific time intervals, where generative performance benefits from finer updates achieved by using smaller time steps. Building on this observation, we extended our analysis to the training stage, exploring two main questions:
@@ B.2 Interaction between Sample and Train Distortions
Appendix C Theoretical Results
In this section, we provide the proofs of the different theoretical results of the paper. We first provide the results that are domain agnostic, i.e., that hold for any data modality that lies in discrete state-spaces, and then the graph specific ones.
C.1 Domain Agnostic Theoretical Results
Here, we first provide the proof of Theorem 2 in Sec. C.1.1, and then the proof of Theorem 3 in Sec. C.1.2.
C.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
We start by introducing two important concepts, which will reveal important for the proof of the intended result.
@@ C.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
As exposed in Sec. 3.1, the marginal distribution and the rate matrix of a CTMC are related by the Kolmogorov equation:
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;R^{d}_{t}(z^{1:D}_{t},z^{d}_{t+\mathrm{d}t})$
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus d}_{t},z^{1:D\setminus d}_{t+\mathrm{d}t})\;\mathbb{E}_{p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}\left[R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z_{1}^{d})\right].$ (13)
In the first equality, ${1:D\setminus d}$ refers to all dimensions except $d$ and the $\delta$ term restricts contributions to rate matrices that account for at most one dimension transitioning at a time, since the probability of two or more independently noised dimensions transitioning simultaneously is zero under a continuous time framework (Campbell et al., 2022; 2024). In the second equality, the unconditional rate matrix is retrieved by taking the expectation over the $z_{1}$-conditioned rate matrices. Specifically, $R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z_{1}^{d})$ denotes the univariate rate matrix corresponding to dimension $d$ (see Eq. 3)
@@
$|R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})|^{2}\leq C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right],$
@@
  • $C_{0}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{d}}\}\sum_{z^{d}_{1}\in\mathcal{Z}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$;
  • $C_{1}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{d}}\}/p_{1}(z^{1:D}_{1})$;
with $C_{z^{d}}=\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\underset{z^{d}_{1}\in\mathcal{Z}}{\sup}\{R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})\}$.
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;R^{d}_{t}(z^{1:D}_{t},z^{d}_{t+\mathrm{d}t})$
$=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\mathbb{E}_{p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}\left[R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})\right]$
$=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\sum_{z^{d}_{1}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})$
@@
$=\left|\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+dt})\;\sum_{z^{d}_{1}}[R^{*d}_{t}(z^{d}_{t},z^{d}_{t+dt}|z^{d}_{1})\left(p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})-p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right)]\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+dt})\;\left|\sum_{z^{d}_{1}}\left[R^{*d}_{t}(z^{d}_{t},z^{d}_{t+dt}|z^{d}_{1})\left(p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})-p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right)\right]\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+dt})\;\underset{z^{d}_{1}}{\sup}\{R^{*d}_{t}(z^{d}_{t},z^{d}_{t+dt}|z^{d}_{1})\}\sum_{z^{d}_{1}}\left|p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})-p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right|$
$=\sum^{D}_{d=1}2\,C_{z^{d}}\,\|p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})-p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\|_{\text{TV}}$ (16)
$\leq\sum^{D}_{d=1}C_{z^{d}}\sqrt{2D_{\text{KL}}\left(p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\,\|\,p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right)}$ (17)
$=\sum^{D}_{d=1}\sqrt{2\;C^{2}_{z^{d}}\sum_{z^{d}_{1}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log\frac{p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}{p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}}.$
In Eq. 16, we use the definition of TV distance as defined in Eq. 14 and Eq. 17 results from direct application of Pinsker’s inequality. Now, we change the ordering of the sum and of the square root through the Cauchy-Schwarz inequality:

### v3

@@ B.3 Detailed Balance, Prior Incorporation, and Stochasticity
still satisfy the Kolmogorov equation. The detailed balance condition ensures that the outflow, $p_{t|1}(z_{t}|z_{1})R^{\text{DB}}_{t}(z_{t},z_{t+\text{d}t}|z_{1})$, and inflow, $p_{t|1}(z_{t+\text{d}t}|z_{1})R^{\text{DB}}_{t}(z_{t+\text{d}t},z_{t}|z_{1})$, of probability mass to any given state are perfectly balanced. Under these conditions, this additive component’s contribution to the Kolmogorov equation becomes null (similar to the target guidance, as shown in the proof of of Lemma 10, in Sec. B.2).
@@
B.4 Hyperparameter Optimization Pipeline
A significant advantage of flow matching methods is their inherently greater flexibility in the sampling process compared to diffusion models, as they are more disentangled from the training stage. Each of the proposed optimization strategies exposed in Sec. 3.2 expands the search space for optimal performance. However, conducting a full grid search across all those methodologies is impractical for the computational resources available. To address this challenge, our sampling optimization pipeline consists of, for each of the proposed optimization strategies, all hyperparameters are held constant at their default values except for the parameter controlling the chosen strategy, over which we perform a sweep. The optimal values obtained for each strategy are combined to form the final configuration. In Tab. 5, we present the final hyperparameter values obtained for each dataset. This pipeline is sufficient to achieve state-of-the-art performance, which reinforces the expressivity of DeFoG. We expect to achieve even better results if a more comprehensive search of the hyperparameter space was carried out.
To better mode detailedly illustrate the influence of each sampling optimization, we show in Figures 10 and 12 in present the impact of varying parameter values across the synthetic datasets Figure 10 and molecular datasets in Figure 12 used in this work.
@@ B.4 Hyperparameter Optimization Pipeline
Figure 12: Sampling efficiency improvement over all molecular datasets. The Guacamol and MOSES datasets are evaluated with 2,000 samples, and validity and novelty are computed using local implementations for efficiency instead of the original benchmarks. For the Guacamol dataset, we apply validity metrics that account for charged molecules, reflecting the dataset’s intrinsic characteristic of containing a significant number of such molecules.
While Figures 10 and 12 are highly condensed, we provide a more fine-grained version that specifically illustrates the influence of the hyperparameters $\eta$ and $\omega$. This version highlights their impact when generating with the full number of steps (500 and 1000 for molecular and synthetic data, respectively) and with 50 steps. As emphasized in Figures 13 and 14, the influence of these hyperparameters varies across datasets and exhibits distinct behaviors depending on the number of steps used.
@@
B.5 Understanding the Sampling Dynamics of $R^{*}_{t}$
In this section, we aim to provide deeper intuition into the sampling dynamics imposed by the design of $R^{*}_{t}$, as proposed by (Campbell et al., 2024). The explicit formulation of $R^{*}_{t}$ can be found in Eq. 3. Notably, the denominator in the expression serves as a normalizing factor, meaning the dynamics of each sampling step are primarily influenced by the values in the numerator. Specifically, we observe the following relationship:
@@ B.5 Understanding the Sampling Dynamics of $R^{*}_{t}$
Table 3: Values of the numerator of $R^{*}_{t}$ for different $z_{t}$ and $z_{t+\mathrm{d}t}$.
Condition Numerator of $R^{*}(x_{t},j|z_{1})$ Intuition
@@
From the first two lines of Tab. 3, we observe that once the system reaches the predicted state $z_{1}$, it remains there. If not, $R^{*}_{t}$ only encourages transitions to other states under two conditions: either the target state is $z_{1}$ (third line), or the corresponding entries in the initial distribution for potential next states have smaller values than the current state (fourth line). As a result, the sampling dynamics are heavily influenced by the initial distribution, as discussed further in Sec. C.1.
@@
For instance, with the masking distribution, the fourth line facilitates transitions to states other than the virtual “mask” state, whereas for the uniform distribution, no transitions are allowed. For the marginal distribution, transitions are directed toward less likely states. Note that while these behaviors hold when the rate matrix consists solely of $R^{*}_{t}$, additional transitions can be introduced through $R^{\mathrm{DB}}_{t}$ (as detailed in Sec. B.3) or by applying target guidance (see Sec. B.2).
B.6 Performance Improvement for Undertrained Models
@@
In this section, we present the performance of a model trained on the QM9 dataset and the Planar dataset using only $30\%$ of the epochs compared to the final model being reported. We employ the same hyperparameters as in Tab. 8 and Tab. 1 for the sampling setup, as reported in Tab. 5.
@@
Compared to fully trained models, our model achieves 99.0 Validity (vs 99.3) and 96.4 Uniqueness (vs 96.3) on the QM9 dataset. For the Planar dataset, it attains 95.5 Validity (vs 99.5) and an average ratio of 1.4 (vs 1.6). These results demonstrate that, even with significantly fewer training epochs, the model maintains competitive performance under a well-designed sampling procedure, although extended training can still further improve performance. Notably, all metrics surpass the discrete-time diffusion benchmark DiGress (Vignac et al., 2022). As a result, the optimization in the sampling stage proves particularly beneficial when computational resources are limited, by enhancing the performance of an undertrained model.
Appendix C Train Optimization
@@
In this section, we provide a more detailed analysis of the influence of the various training optimization strategies introduced in Sec. 3. In Sec. C.1, we empirically demonstrate the impact of selecting different initial distributions on performance, while in Sec. C.2, we examine the interaction between training and sampling optimization.
C.1 Initial Distributions
@@
Under DeFoG’s framework, the noising process for each dimension is modeled as a linear interpolation between the clean data distribution (the one-hot representation of the current state) and an initial distribution, $p_{0}$. As such, it is intuitive that different initial distributions result in varying performances, depending on the denoising dynamics they induce. In particular, they have a direct impact on the sampling dynamics through $R^{*}_{t}$ (see Sec. B.5) and may also pose tasks of varying difficulty for the graph transformer. In this paper, we explore four distinct initial distributions [^2] Recall that $Z$ represents the cardinality of the state space, and $\Delta^{Z}$ the associated probability simplex.:
@@
Uniform: $p_{0}=\left[\frac{1}{Z},\frac{1}{Z},\ldots,\frac{1}{Z}\right]\in\Delta^{Z}$. Here, the probability mass is uniformly distributed across all states, as proposed by (Campbell et al., 2024).
@@
Masking: $p_{0}=[0,0,\ldots,0,1]\in\Delta^{Z+1}$. In this setting, all the probability mass collapses into a new “mask” state at $t=0$, as introduced by (Campbell et al., 2024).
@@
Marginal: $p_{0}=[m_{1},m_{2},\ldots,m_{Z}]\in\Delta^{Z}$, where $m_{i}$ denotes the marginal probability of the $i$-th state in the dataset. This approach is widely used in state-of-the-art graph generation models (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024).
@@
Absorbing: $p_{0}=[0,\ldots,1,\ldots,0]\in\Delta^{Z}$, representing a one-hot encoding of the most common state (akin to applying an $\operatorname{argmax}$ operator to the marginal initial distribution).
@@
We apply each of these initial distributions to both node and edges, with the corresponding dimensionalities. In Figure 19, we present the training curves for each initial distribution for three different datasets.
@@
We observe that the marginal distribution consistently achieves at least as good performance as the other initial distributions. This is supported by two key observations in previous work: first, Vignac et al. (2022) demonstrate that, for any given distribution, the closest distribution within the aforementioned class of factorizable initial distributions is the marginal one, thus illustrating its optimality as a prior. Second, the marginal initial distribution preserves the dataset’s marginal properties throughout the noising process, maintaining graph-theoretical characteristics like sparsity (Qin et al., 2023). We conjecture that this fact facilitates the denoising task for the graph transformer. This reinforces its use as the default initial distribution for DeFoG. The only dataset where marginal was surpassed was the SBM dataset, which we attribute to its inherently different nature (stochastic, instead of deterministic). In this case, the absorbing distribution emerged as the best-performing choice. Interestingly, the absorbing distribution tends to converge faster across datasets.
@@
C.2 Interaction between Sample and Train Distortions
From Sec. B.4, we observe that time distortions applied during the sampling stage can significantly affect performance. This suggests that graph discrete flow models do not behave evenly across time and are more sensitive to specific time intervals, where generative performance benefits from finer updates achieved by using smaller time steps. Building on this observation, we extended our analysis to the training stage, exploring two main questions:
@@ C.2 Interaction between Sample and Train Distortions
Appendix D Theoretical Results
In this section, we provide the proofs of the different theoretical results of the paper. First, we provide results that are domain agnostic, i.e., that hold for general discrete data, and then instantiate them for the specific case of graphs, yielding Corollaries 1 and 2. Then, we proceed to the permutation invariance/equivariance guarantees of DeFoG, and some remarks on the expressivity of the RRWP additional features.
D.1 Theoretical Results on the Generative Framework
In Sec. D.1.1, we provide a proof of the bounded estimation error of the rate matrix for general discrete data, and instantiate it for graphs, yielding Corollary 1. Then, in Sec. D.1.2, we prove the bounded deviation of the generated distribution for general discrete data as well, and instantiate it for graphs, yielding Corollary 2.
Importantly, for general discrete data, we jointly model $D$ discrete variables, $(z^{(1)},\ldots,z^{(D)})=z^{1:D}\in\mathcal{Z}^{D}$.
D.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
We start by introducing two important concepts, which will reveal useful for the proof of the intended result.
@@ D.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
As exposed in Sec. 2, the marginal distribution and the rate matrix of a CTMC are related by the Kolmogorov equation:
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;R^{(d)}_{t}(z^{1:D}_{t},z^{(d)}_{t+\mathrm{d}t})$ (10)
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\mathbb{E}_{p_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}\left[R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z_{1}^{(d)})\right].$ (11)
In the first equality, ${1:D\setminus(d)}$ refers to all dimensions except $d$ and the $\delta$ term restricts contributions to rate matrices that account for at most one dimension transitioning at a time, since the probability of two or more independently noised dimensions transitioning simultaneously is zero under a continuous time framework (Campbell et al., 2022, 2024). In the second equality, the unconditional rate matrix is retrieved by taking the expectation over the $z_{1}$-conditioned rate matrices. Specifically, $R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z_{1}^{(d)})$ denotes the univariate rate matrix corresponding to dimension $d$ (see Eq. 3)
@@
$|R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})|^{2}\leq C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{{t|1}}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right],$
@@
  • $C_{0}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{(d)}}\}\sum_{z^{(d)}_{1}\in\mathcal{Z}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$;
  • $C_{1}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{(d)}}\}/p_{1}(z^{1:D}_{1})$;
with $C_{z^{(d)}}=\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\underset{z^{(d)}_{1}\in\mathcal{Z}}{\sup}\{R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})\}$.
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;R^{(d)}_{t}(z^{1:D}_{t},z^{(d)}_{t+\mathrm{d}t})$
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\mathbb{E}_{p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}\left[R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})\right]$
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\sum_{z^{(d)}_{1}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})$
@@
$=\left|\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+dt})\;\sum_{z^{(d)}_{1}}[R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+dt}|z^{(d)}_{1})\left(p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})-p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right)]\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+dt})\;\left|\sum_{z^{(d)}_{1}}\left[R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+dt}|z^{(d)}_{1})\left(p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})-p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right)\right]\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+dt})\;\underset{z^{(d)}_{1}}{\sup}\{R^{(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+dt}|z^{(d)}_{1})\}\sum_{z^{(d)}_{1}}\left|p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})-p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right|$
$=\sum^{D}_{d=1}2\,C_{z^{(d)}}\,\|p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})-p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\|_{\text{TV}}$ (14)
$\leq\sum^{D}_{d=1}C_{z^{(d)}}\sqrt{2D_{\text{KL}}\left(p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\,\|\,p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right)}$ (15)
$=\sum^{D}_{d=1}\sqrt{2\;C^{2}_{z^{(d)}}\sum_{z^{(d)}_{1}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log\frac{p^{(d)}(_{1|t}z^{(d)}_{1}|z^{1:D}_{t})}{p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}}.$